# A Context for Manifold Calculus

### v2

@@ Abstract.
1. Introduction
Let $M$ be a smooth $n$-manifold and let $\operatorname{Open}\left(M\right)$ be the poset of open sets of $M$. The idea of Weiss’s manifold calculus is to study space-valued presheaves $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ on $M$ by constructing a tower of presheaves
$F\to\cdots\to T_{k}F\to\cdots\to T_{1}F\to T_{0}F$
such that each $T_{k}F$ is the best “polynomial functor” of degree $\leq k$ which approximates $F$. Borrowing an analogy from calculus, the tower is called the Taylor tower of $F$ and the functors $T_{k}F$ the polynomial approximations of $F$.
The Taylor tower is determined by how we define what it means for a functor to be polynomial, so we must make this explicit. There is a heuristic in homotopy theory that homology behaves like linear (or affine) functions. A homology theory is more or less characterized by the excision axiom, so we make the following definition:
@@ 1. Introduction
$FU\to\underset{\emptyset\subsetneq S\subset\{0,\dots,k\}}{\operatorname{holim}}F\left(U\setminus\bigcup_{i\in S}A_{i}\right)$
@@
For example, $1$-excisivity is equivalent to the condition that, for every open set $U\subset M$ and every pair of disjoint closed sets $A_{0},A_{1}\subset U$, the square
${{F(U)}}$${{F(U\setminus A_{1})}}$${{F(U\setminus A_{0})}}$${{F(U\setminus(A_{0}\cup A_{1}))}}$
@@
  • Let $U\subset V$ be an inclusion of open sets of $M$ which is a smooth isotopy equivalence. Then the map $FV\to FU$ is a homotopy equivalence.
@@
Suppose now that $F$ takes values in the category of pointed spaces and that we have constructed its polynomial approximations. In good cases, we can expect that the analysis of $F$ reduces, up to extension problem, to that of the fibers $L_{k}F=\mathrm{hofib}\left(T_{k}F\to T_{k-1}F\right)$. The functor $L_{k}F$ has the special property that it is homogeneous: It is polynomial of degree $\leq k$, and its $\left(k-1\right)$th polynomial approximation vanishes. This leads to the following question:
Question 1.6.
@@
And of course, we want to know:
Question 1.7.
Is manifold calculus actually useful?
Numerous studies suggest that the answer to Question 1.7 is an unequivocal yes: The idea of manifold calculus is a very, very good idea. Rather than making a long list of applications, we refer the reader to Munson’s survey [Mun10]. As for the remaining questions, Weiss gave complete answers in [Wei99], which we now review as two separate theorems (Theorems 1.8 and 1.9).
The first theorem answers Questions 1.4 and 1.5. We will let $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)$ denote the subset of $\operatorname{Open}\left(M\right)$ consisting of the elements that are diffeomorphic to $\mathbb{R}^{n}\times S$, where $S$ is a finite set of cardinality at most $k$.
@@
[Wei99, Lemma 3.8, Theorem 4.1, Theorem 5.1] Let $G:\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ be a functor which carries isotopy equivalences to homotopy equivalences. The homotopy right Kan extension of $G$ along the inclusion $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\hookrightarrow\operatorname{Open}\left(M\right)^{\mathrm{op}}$ is good and $k$-excisive. Moreover, every good polynomial functor of degree $\leq k$ arises in this way.
@@
And the second theorem addresses Question 1.6:
@@
[Wei99, Theorem 8.5] Homogeneous functors of degree $k$ taking values in pointed spaces admit a classification in terms of fibrations over $B_{k}\left(M\right)$ equipped with sections, where $B_{k}\left(M\right)$ denotes the space of unordered configurations of $k$ points in $M$.
Now notice that everything we have discussed so far makes sense for presheaves taking values in categories other than that of spaces. Moreover, it seems quite natural to consider this generalization. For example, if one is interested in (co)homology, then it is not hard to imagine that calculi of presheaves of (co)chain complexes or spectra would be very useful. As such, numerous studies consider such generalized manifold calculus. But none of them seem to address Questions 1.4, 1.5, and 1.6 in this generalized setting. We will fix this in this paper.
Here is what we will do. We will generalize Theorems 1.8 and 1.9 to presheaves on $M$ taking values in an arbitrary $\infty$-category $\mathcal{C}$ with small limits. It is straightforward to state a generalization of Theorem 1.8: The definitions of good functors, polynomial functors, polynomial approximations, and homogeneous functors directly carry over to $\mathcal{C}$-valued presheaves; we just have to replace $\operatorname{Open}\left(M\right)$ by its nerve $N\left(\operatorname{Open}\left(M\right)\right)$ and homotopy limits by $\infty$-categorical limits. So a generalization of Theorem 1.8 will be as follows.
@@
$\operatorname{Fun}_{{\rm istp}}\left(N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}},\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right),$
@@
  • $\operatorname{Fun}_{{\rm istp}}\left(N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}},\mathcal{C}\right)$ denotes the $\infty$-category of functors $N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$ which carries smooth isotopy equivalences to equivalences; and
  • $\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of $k$-excisive good functors $N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$.
In particular, every good functor $F:N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$ admits a $k$th polynomial approximation, which is the right Kan extension of $F|N\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\right)^{\mathrm{op}}$.
To generalize Theorem 1.9, notice that the datum of a fibration over $B_{k}\left(M\right)$ equipped with a distinguished section is equivalent (via the $\infty$-categorical Grothendieck construction [Lur09, 2.2.1.2]) to that of a functor $\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{S}_{\ast}$, where $\mathcal{S}_{\ast}$ denotes the $\infty$-category of pointed spaces. With this in mind, we generalize Theorem 1.9 as follows:
Theorem 1.11 (Corollary 4.5).
@@
where $\operatorname{Homog}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of good functors $N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}$ that are of homogeneous of degree $k$.
@@
Notice that most of the statements of Theorems 1.10 and 1.11 make sense even when $M$ does not have a smooth structure. Because of this, we will work mostly with topological manifolds in the main body of the paper, indicating how various statements will be affected in the presence of smooth structures here and there.
@@
In [STS19], Songhafouo Tsopméné and Stanley considered the classification of homogeneous functors in the case where $\mathcal{C}$ is the nerve of an ordinary category. Their result is a special case of Theorem 1.11, although our proof method is quite different from theirs.
At first glance, one may think that proofs of Theorem 1.10 and 1.11 are straightforward, using a minor modification of Weiss’s argument. However, Weiss’s original argument does not seem to generalize to the case at hand, for it relies critically on concrete constructions of spaces. We will get around this problem by using a deep theorem on localizations with respect to isotopy equivalences, due to Ayala and Francis [AF15, Proposition 2.19] [^2] Since the original proof of [AF15, Proposition 2.19] is missing some important justifications, we will record a detailed proof in Subsection 2.3. The author learned the proof from David Ayala and is grateful to him for explaining the proof..
Theorem 1.10 and 1.11 have a wide array of applications outside what has traditionally been regarded as part of manifold calculus. For example, (part of) the theory of factorization homology [AF20] deals with functors of the form
@@
where $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denotes the $\infty$-category of smooth $n$-manifolds and smooth embeddings. By precomposing the functor $N\left(\operatorname{Open}\left(M\right)\right)\to\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, we obtain a $\mathcal{C}^{\mathrm{op}}$-valued precosheaf $F:N\left(\operatorname{Open}\left(M\right)\right)^{\mathrm{op}}\to\mathcal{C}^{\mathrm{op}}$. An observation of Ayala and Francis [AF20, Corollary 2.2.24] (which will be reproved in Subsection 3.4) essentially asserts that the functor $F$ is the limit of its Taylor tower, in the sense that the map $F\to\operatorname{lim}_{k}T_{k}F$ is an equivalence. This means that invariants of manifolds realized by factorization homology, which include very interesting class of invariants [AF20], can be approximated by their Taylor approximations. As another example, we will apply Theorem 1.11 in our future work to a functor taking values in the $\infty$-category $\mathcal{O}\mathsf{p}$ of $\infty$-operads. We will see that Theorem 1.11 is just the right tool to establish that this functor enjoys a “local-to-global” property. In light of the ubiquity of pre(co)sheaves on manifolds, we expect that there will be many other applications.
@@
In Section 2, we collect miscellaneous results that will be used in later sections. The statements of the results in later sections can be understood without reading this section, so we encourage readers to read this section only when the need arises. Section 3 concerns the existence and classification of polynomial functors, as well as the convergence of Taylor towers. Section 4 classifies homogeneous functors. In Section 5, we prove a generalization of the results in Section 3 for context-free manifold calculus. Appendix A contains a discussion of the context-free analog of the localization theorem of Ayala and Francis. The results in the appendix will not be used elsewhere, but it has some independent interest.
@@ Acknowledgment
I appreciate David Ayala for kindly sharing the proof of Theorem 2.24, which plays a central role in this paper. Alexander Kupers generously answered my question on MathOverflow [sku]. His answer is presented as Proposition 2.7. I am also indebted to Daisuke Kishimoto and Mitsunobu Tsutaya for their constant support and encouragement.
@@
  • We say that a simplicial category $\mathcal{C}$ is locally Kan if its hom-simplicial sets are Kan complexes. A full simplicial subcategory $\mathcal{C}^{0}\subset\mathcal{C}$ is homotopically replete if $X\to Y$ is a homotopy equivalence of $\mathcal{C}$ and $X\in\mathcal{C}^{0}$, then $Y\in\mathcal{C}^{0}$.
  • A functor $\mathcal{C}\to\mathcal{D}$ of simplicial categories is called a weak equivalence if it induces a categorical equivalence $\pi_{0}\left(\mathcal{C}\right)\to\pi_{0}\left(\mathcal{D}\right)$ and a weak homotopy equivalence $\mathcal{C}\left(X,Y\right)\to\mathcal{D}\left(fX,fY\right)$ for every pair of objects $X,Y\in\mathcal{C}$.
@@
    • We say that a functor $f:\mathcal{C}\to\mathcal{D}$ of $\infty$-categories is a localization with respect to a set $S$ of morphisms of $\mathcal{C}$ if it is initial among functors inverting the morphisms in $S$. (See Subsection 2.3 for a precise definition.) This definition is more general than the definition of [Lur09].
@@
  • Let $n\geq 0$. An $n$-manifold, or a manifold of dimension $n$, will always mean a topological manifold without boundary, i.e., a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$. If $n\geq 1$, we define an $n$-manifold with boundary to be a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$ or $[0,\infty)\times\mathbb{R}^{n-1}$. In particular, not every $n$-manifold with boundary is an $n$-manifold.
@@
  • Let $n\geq 0$. We let $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the homotopy coherent nerve of the topological category of smooth $n$-manifolds and smooth embeddings, whose mapping spaces $\operatorname{Emb}_{\mathrm{sm}}\left(-,-\right)$ are topologized by the weak topology (also called the compact-open $C^{\infty}$ topology) of [Hir76, Chapter 2]. We let $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}\subset\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the full subcategory spanned by the smooth manifolds diffeomorphic to a finite disjoint union of $\mathbb{R}^{n}$ (with the standard smooth structure). Equivalences of $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ will be called smooth isotopy equivalences. The $\infty$-categories $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, $\mathsf{Mfld}_{\mathrm{sm},n}$, $\mathsf{Disk}_{\mathrm{sm},n}$, and $\mathsf{Disk}_{\mathrm{sm},n}^{\leq k}$ are defined as in the previous point.
  • Let $n,k\geq 0$. If $M$ is an $n$-manifold, we will write $\operatorname{Disj}\left(M\right)\subset\operatorname{Open}\left(M\right)$ for the full subcategory spanned by the open sets homeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$, and let $\operatorname{Disj}^{\leq k}\left(M\right)$ denote its full subposet spanned by the open sets with at most $k$ components. If $M$ is smooth, we let $\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\subset\operatorname{Disj}\left(M\right)$ denote the full subcategory spanned by open sets diffeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$ (with the standard smooth structure), and write $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)=\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\cap\operatorname{Disj}^{\leq k}\left(M\right)$.
@@
2. Preliminaries
In this section, we develop machineries we will use in later sections. Subsection 2.1 deals with germ bundles, which is an analog of tangent bundles in the topological setting. In Subsection 2.2, we consider the relation between Kan extensions and colimit diagrams, using Lurie’s path fibrations. In Subsection 2.3, we review Ayala and Francis’s localization theorems for isotopy equivalences. Finally, in Subsection 2.4, we consider the problem of zero extensions of functors.
The first three subsections are written in a linear order. The final subsection is independent of the other subsections.
2.1. Germs

### v3

@@ Abstract.
Introduction
Let $M$ be a smooth $n$-manifold, and let $\operatorname{Open}\left(M\right)$ be the poset of open sets of $M$. The idea of Weiss’s manifold calculus is to study space-valued presheaves $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ on $M$ by constructing a tower of presheaves
$F\to\cdots\to T_{k}F\to\cdots\to T_{1}F\to T_{0}F.$
Each $T_{k}F$ is the best “polynomial functor” of degree $\leq k$ that approximates $F$. Borrowing an analogy from calculus, the tower is called the Taylor tower of $F$ and the functors $T_{k}F$ the polynomial approximations of $F$.
The Taylor tower is determined by the definition of polynomials, so we must make this explicit. There is a heuristic in homotopy theory that homology behaves like linear (or affine) functions. A homology theory is more or less characterized by the excision axiom, so we make the following definition:
@@ Introduction
$F\left(U\right)\to\underset{\emptyset\subsetneq S\subset\{0,\dots,k\}}{\operatorname{holim}}F\left(U\setminus\bigcup_{i\in S}A_{i}\right)$
@@
For example, $1$-excisivity is equivalent to the following condition: For every open set $U\subset M$ and every pair of disjoint closed sets $A_{0},A_{1}\subset U$, the square
[…]
@@
  • Let $U\subset V$ be an inclusion of open sets of $M$ which is a smooth isotopy equivalence. Then the map $F\left(V\right)\to F\left(U\right)$ is a homotopy equivalence.
@@
Suppose now that $F$ takes values in the category of pointed spaces and that we have constructed its polynomial approximations. In good cases, the analysis of $F$ reduces, up to extension problems, to studying the fibers $L_{k}F=\mathrm{hofib}\left(T_{k}F\to T_{k-1}F\right)$. The functor $L_{k}F$ has the special property that it is homogeneous: It is polynomial of degree $\leq k$, and its $\left(k-1\right)$th polynomial approximation vanishes. This leads to the following question:
Question 0.6.
@@
In [Wei99], Weiss gave complete answers to these questions. We now review his answers as two separate theorems (Theorems 0.7 and 0.8).
The first theorem answers Questions 0.4 and 0.5. We will let $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)$ denote the subset of $\operatorname{Open}\left(M\right)$ consisting of the elements that are diffeomorphic to $\mathbb{R}^{n}\times S$, where $S$ is a finite set of cardinality at most $k$.
@@
[Wei99, Lemma 3.8, Theorem 4.1, Theorem 5.1] Let $G:\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\to\mathsf{Spaces}$ be a functor carrying isotopy equivalences to homotopy equivalences. The homotopy right Kan extension of $G$ along the inclusion $\operatorname{Disj}_{{\rm sm}}^{\leq k}\left(M\right)^{\mathrm{op}}\hookrightarrow\operatorname{Open}\left(M\right)^{\mathrm{op}}$ is good and $k$-excisive. Moreover, every good polynomial functor of degree $\leq k$ arises in this way.
@@
And the second theorem addresses Question 0.6:
@@
[Wei99, Theorem 8.5] Homogeneous functors of degree $k$ taking values in pointed spaces admit a classification in terms of fibrations over $B_{k}\left(M\right)$ equipped with sections near the “fat diagonal”, where $B_{k}\left(M\right)$ denotes the space of unordered configurations of $k$ points in $M$.
Most of the things (perhaps except for Theorem 0.8) we have discussed so far make sense for presheaves taking values in categories other than that of spaces. Moreover, it seems quite natural to consider this generalization. For example, if one is interested in (co)homology, then it is not hard to imagine that presheaves of (co)chain complexes or spectra would be very useful. As such, numerous studies consider such generalized manifold calculus; see [Wei04, RW14, HK23, SAS18, TS25] for instance. In this paper, we will answer Questions 0.4, 0.5, and 0.6 in the generalized setting, and aim to set a foundation of generalized manifold calculus with arbitrary targets.
Here is what we will do. We will generalize Theorems 0.7 and 0.8 to presheaves on $M$ taking values in an arbitrary $\infty$-category $\mathcal{C}$ with small limits. It is straightforward to state a generalization of Theorem 0.7: The definitions of good functors, polynomial functors, polynomial approximations, and homogeneous functors directly carry over to $\mathcal{C}$-valued presheaves; we just have to replace homotopy limits by $\infty$-categorical limits. So a generalization of Theorem 0.7 will be as follows.
@@
$\operatorname{Fun}_{{\rm istp}}\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)\xrightarrow{\simeq}\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right),$
@@
  • $\operatorname{Fun}_{{\rm istp}}\left(\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)^{\mathrm{op}},\mathcal{C}\right)$ denotes the $\infty$-category of functors $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)\to\mathcal{C}$ carrying smooth isotopy equivalences to equivalences; and
  • $\operatorname{Exc}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of $k$-excisive good functors $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$.
In particular, every good functor $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ admits a $k$th polynomial approximation, which is the right Kan extension of $F|\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)^{\mathrm{op}}$.
To generalize Theorem 0.8, recall that a fibration over $B_{k}\left(M\right)$ is equivalent (via the $\infty$-categorical Grothendieck construction [Lur09, 2.2.1.2]) to a functor $\operatorname{Sing}B_{k}\left(M\right)\to\mathcal{S}$, where $\mathcal{S}$ denotes the $\infty$-category of spaces. With this in mind, we generalize Theorem 0.8 as follows:
Theorem 0.10 (Corollary 3.5, Theorem 3.16).
@@
where $\operatorname{Homog}_{{\rm good}}^{k}\left(M^{\mathrm{op}};\mathcal{C}\right)$ denotes the $\infty$-category of good functors $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}$ that are of homogeneous of degree $k$.
@@
Theorem 0.10 says nothing about the fat diagonal, but a more precise statement involving the fat diagonal statement will follow. (See Theorem 3.16.) In particular, when applied to the case where $\mathcal{C}$ is the $\infty$-category $\mathcal{S}_{\ast}$ of pointed spaces, the theorem recovers Weiss’s classification of homogeneous functors.
Remark 0.12.
Most of the statements of Theorems 0.9 and 0.10 make sense even when $M$ does not have a smooth structure. Because of this, we will work mostly with topological manifolds, explaining how various statements will be affected in the presence of smooth and PL structures here and there.
@@
Our paper is thematically very similar to the works of Songhafouo Tsopméné, Arnaud, and Stanley[SAS18, SAS19, TS25], in which the authors ask the extent to which manifold calculus can be developed for presheaves taking values in categories other than spaces. (They mainly work in the setting of model categories and employ very different technique than ours.) Our theorems will recover the results in their paper and solve their conjecture. More specifically:
  • By applying Theorem 0.9 and 0.10 to the case where $\mathcal{C}$ is the underlying $\infty$-category of a simplicial model category, we recover the main results of [SAS18].
  • By applying Theorem 0.10 to the case where $\mathcal{C}$ is an ordinary category, we recover the main result of [SAS19].
  • By applying Theorem 0.10 to the case where $\mathcal{C}$ is the underlying $\infty$-category of a simplicial model category $\mathbf{C}$, we recover the main result and a conjecture of [TS25]: For each object $X\in\mathbf{C}$, the homotopy classes of maps $B_{k}\left(M\right)\to B\mathrm{hAut}\left(X\right)$ are in bijection with the weak equivalence class good functors $\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathbf{C}$ that are of homogeneous of degree $k$ and carrying disjoint union of $k$ open balls to objects weakly equivalent to $X$.
At first glance, one might expect that proofs of Theorem 0.9 and 0.10 are straightforward, requiring only a minor modification of Weiss’s argument. While this is true to some extent, several parts of Weiss’s original proof rely critically on concrete constructions of spaces, which do not generalize easily. Moreover, our theorems is slightly stronger than Weiss’s results, in that it gives an equivalence of $\infty$-categories, and this added generality introduces further complications. To overcome these challenges, we develop and apply more sophisticated $\infty$-categorical techniques.
Theorem 0.9 and 0.10 have a wide array of applications outside what has traditionally been regarded as part of manifold calculus. For example, (part of) the theory of factorization homology [AF20] deals with functors of the form
@@
where $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denotes the $\infty$-category of smooth $n$-manifolds and smooth embeddings. By precomposing the functor $\operatorname{Open}\left(M\right)\to\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$, we obtain a $\mathcal{C}^{\mathrm{op}}$-valued precosheaf $F:\operatorname{Open}\left(M\right)^{\mathrm{op}}\to\mathcal{C}^{\mathrm{op}}$. An observation of Ayala and Francis [AF20, Corollary 2.2.24] (which will be reproved in Subsection 2) essentially asserts that the functor $F$ is the limit of its Taylor tower, in the sense that the map $F\to\operatorname{lim}_{k}T_{k}F$ is an equivalence. This means that invariants of manifolds realized by factorization homology, which include a very interesting class of invariants [AF20], can be approximated by their Taylor approximations. It is crucial here to have a flexibility in the target category, which this paper offers, because invariants can arise in many different forms, such as spectra, chain complexes, or even $\infty$-operads [Ara24b].
@@
This paper consists of 5 sections. Section 1 concerns the existence and classification of polynomial functors. In Section 2, we construct and consider the convergence problem of Taylor towers. Section 3 classifies homogeneous functors. In Section 4, we prove a generalization of the results in Sections 1 and 2 for context-free manifold calculus. Section 5, discusses how these ideas extend to manifolds with boundary
The paper is also accompanied by an appendix, consisting of three sections. The purpose of the appendix is to record miscellaneous results used in the main body of the paper, so it should be consulted only when the need arises.
@@ Acknowledgment
I thank Michael Weiss for carefully explaining some of the key ideas of his papers, David Ayala for kindly sharing the proof of Theorem A.1, and Alexander Kupers for generously answering my question on MathOverflow [sku]. I am also indebted to Daisuke Kishimoto and Mitsunobu Tsutaya for their constant support and encouragement. Last, but not least, I express my sincerest gratitude to the referee for carefully reading the manuscript and making valuable feedback, which has significantly improved this paper.
@@
    • We say that a functor $f:\mathcal{C}\to\mathcal{D}$ of $\infty$-categories is a localization with respect to a set $S$ of morphisms of $\mathcal{C}$ if it is initial among functors inverting the morphisms in $S$. (See Subsection A.1 for a precise definition.) This definition is more general than the definition in [Lur09].
    • Except in Appendix A, we will not distinguish between ordinary categories and their nerves.
@@
  • Let $n\geq 0$. An $n$-manifold, or a manifold of dimension $n$, will always mean a topological manifold without boundary, i.e., a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$. If $n\geq 1$, we define an $n$-manifold with boundary to be a second countable, Hausdorff topological space which admits an open cover by open sets homeomorphic to $\mathbb{R}^{n}$ or $0,\infty)\times\mathbb{R}^{n-1}$. In particular, not every $n$-manifold with boundary is an $n$-manifold.
@@
  • Let $n\geq 0$. We let $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the homotopy coherent nerve of the topological category of smooth $n$-manifolds and smooth embeddings, whose mapping spaces $\operatorname{Emb}_{\mathrm{sm}}\left(-,-\right)$ are topologized by the weak topology (also called the compact-open $C^{\infty}$ topology) of [Hir76, Chapter 2]. We let $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}\subset\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ denote the full subcategory spanned by the smooth manifolds diffeomorphic to a finite disjoint union of $\mathbb{R}^{n}$ (with the standard smooth structure). Equivalences of $\mathcal{M}\mathsf{fld}_{\mathrm{sm},n}$ will be called smooth isotopy equivalences. The $\infty$-categories $\mathcal{D}\mathsf{isk}_{\mathrm{sm},n}^{\leq k}$, $\mathsf{Mfld}_{\mathrm{sm},n}$, $\mathsf{Disk}_{\mathrm{sm},n}$, and $\mathsf{Disk}_{\mathrm{sm},n}^{\leq k}$ are defined as in the previous point. (We will not define a PL version of these $\infty$-categories because the author is not aware of a good topology on PL embeddings whose paths correspond to PL isotopies.)
  • Let $n,k\geq 0$. If $M$ is an $n$-manifold, we will write $\operatorname{Disj}\left(M\right)\subset\operatorname{Open}\left(M\right)$ for the full subcategory spanned by the open sets homeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$, and let $\operatorname{Disj}^{\leq k}\left(M\right)$ denote its full subposet spanned by the open sets with at most $k$ components. If $M$ is smooth, we let $\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\subset\operatorname{Disj}\left(M\right)$ denote the full subcategory spanned by the open sets diffeomorphic to $\mathbb{R}^{n}\times S$ for some finite set $S$ (with the standard smooth structure), and write $\operatorname{Disj}_{\mathrm{sm}}^{\leq k}\left(M\right)=\operatorname{Disj}_{\mathrm{sm}}\left(M\right)\cap\operatorname{Disj}^{\leq k}\left(M\right)$.
@@
  • For each finite set $S$, we will write $\mathcal{P}\left(S\right)$ for the poset of subsets of $S$, and $\mathcal{P}_{0}\left(S\right)$ for the poset of nonempty subsets of $S$.
1. Polynomial Functors
In this section, we will show that polynomial approximations exist and that they are constructed by Kan extending along $\operatorname{Disj}^{\leq k}\left(M\right)\to\operatorname{Open}\left(M\right)$ (Theorem 1.3).
Warning 1.1.
To avoid writing $\mathrm{op}$” repeatedly, we will work with covariant functors in this section, so that polynomial functors are described in terms of cosheaves and left Kan extensions. However, in later section, we sometimes work with the dual theory of presheaves and right Kan extensions (such as Theorems 2.7 and 3.16). The choice will be dictated by the ease it takes to state and prove the relevant results.
Let us start by stating the main result of this section (Theorem 1.3).